Miles per gallon of a vehicle is a random variable with a uniform distribution from 23 to 47. The probability that a random vehicle gets between 28 and 36 miles per gallon
Answer #1

Solution :

Given that,

a = 23

b = 47

P(c < x < d) = (d - c) / (b - a)

P(28 < x < 36) = (36 - 28) / (47 - 23) = 0.3333

Probability = 0.3333